Person Edison is a 44-year-old autoworker who was laid off from his job 6 months ago. He is frustrated with his inability toNot in the adult population Employed Not in the adult population Not in the labor force UnemployedComplete the right half of the following equation to reflect the labor force participation rate reported by the BLS. Labor FoComplete the right half of the following equation to reflect the unemployment rate reported by the BLS. Unemployment Rate = A
Answer #1

Edison- employed

Hilary- unemployed

Kevin- not in the labor force

Simone- unemployed

Rajiv- not in the labor force

Maria- not in the adult population

Labor force participation rate = labor force *100 / adult population

= 3*100/5= 60%

Unemployment rate = number of unemployed *100 / labor force

= 2*100/3= 66.67%